[Halld-jlab] Hall D Staff Meeting Notes, July 27, 2015
Mark Ito
marki at jlab.org
Tue Jul 28 17:44:23 EDT 2015
Folks,
Find the notes below and at
https://halldweb.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_27,_2015
-- Mark
___________________
Private:Hall D Staff Meeting, July 27, 2015
Present: Mary Jo Bailey, Fernando Barbosa, Eugene Chudakov (chair),
Alexandre Deur, Sergey Furletov, Mark Ito, David Lawrence, Bert Manzlak,
Paul Mattione, Lubomir Pentchev, Elton Smith, Sascha Somov, Justin
Stevens, Simon Taylor, Glenn Young, Beni Zihlmann
Eugene
* PAC results came out
* solenoid review, two weeks ago
o final report is out
o thermostats should be looked at
o lack of cryostability: cause not clear
+ local heating?
+ ice build up? in He volume? vacuum volume?
o He supply system will not see small fluctuations
+ 20% variation might not be noticed
o pressure rise causes rise in resistance
* we need to make a space request for the DIRC to Walt Akers
o Justin: leave boxes at SLAC as long as possible
o when they are moved here, move directly to Hall
o will try to use one-bar-box read-out prototype set-up at SLAC
o will needs to devise tests to make sure boxes made trip to JLab
safely
Mary Jo
* new proposal for hazard issues list
o list reviewed annually
o part of EH&S manual
o old list revised: some items not strictly issues, per se
o rewritten from worker's point of view
o linked to OSP system to ensure appropriate SME review
o methodology: ask questions for each area
o comments to Mary Jo (mbailey at jlab.org)
o proposed roll-out of new list: August 3, 2105
Fernando
* stuff is working
* current work: polarimeter
o new enclosure, power distribution
* MO boards
o ran without firmware in Spring, Ed working on it
o RF clock jitter: several fs, between crates: 150 ps
o Ed J.: perhaps from loss of resolution at rollover of F1-TDC
o Ben: jitter tested out of clock distribution system: 3 ps
o in Hall there are TIs, STs
o Paul and Elton looking at jitter
o 4 crates involved: PS, TOF, tagger, FDC
Eugene, Engineering
* Tim working on acoustic detectors for solenoid
* work with SRF group
* Piezo-electric device attached to metal plate, from there with rods
to coils
* thermal events make no noise, motion of flux jumps might
Eugene, End-of-Fiscal-Year Purchases
* please send requests to Eugene within the week
* we have lots of TI boards
* we may be short of trigger distribution boards
* rapid access for Hall D requires CARM boxes, not cheap
Mark and David, Git Conversion
* Git conversion generally going well
o No major problems
o We are climbing the learning curve
David
* sending a test set-up to UMass for CPP detector tests
o including an Flash-125
* working on Geant4, merging CPP with Richard Jones's work on geometry
Elton
* loan to UMass went quickly perhaps because CPP is an approved experiment
* BCAL back on after power shutdown
o HV channel tripped, on for one week now
o humidity monitors were erratic, replaced sensors, upstream and
downstream about the same now
* FCAL calibrations
o looking at electrons 10% lower energy for clusters, hits are
included from out-of-time
o see radial dependence in constants
* focus shifting to Monte Carlo, looking for limits of hardware
* improvements to firmware for Flash-250s being discussed
o Sasha: document exists for scintillators
Simon
* FCAL clusters including late hits, re-working clusterizer adding
timing cut
* CDC straws distorted, timing distributions different on different
sides of the wire, source of long tails in timing dists
o Mike Staib looking at cosmics, Simon at beam data
o bad straw fraction may be as high as 50%
Sergey
* looking at new FPGA code for F1TDCs
* asked for final version of F125 code from Cody
o at least would like final version as far as DAQ is concerned
Paul
* submitted proposal to JSA for analysis workshop
* scaler gui's written, need test
Mark
* Spring 2015 simulations: data generated
* Data Challenge 3: generating 10 TB of "raw" data
Sascha
* F1-TDC, Ed, read-out failure at high rate, Sergey reproduced, Ed
confirmed
o new version of firmware coming
o FIFO overflow causes FPGA code failure
* students are documenting, pair spectrometer hodoscope, time calibration
* working on proposal for nuclear targets
Alexandre
* looking at beamline data
o no time coincidences between beam event and magnet quench found
Justin
* S&T starts tomorrow, runs for 3 days
o has given practice talks, favorable reaction from Management
* polarized rho's, fits being performed, spin density matrix elements
look do-able
* DIRC: prepare for review, writing TDR documents
Lubomir
* FDC at design resolution with straight tracks
o 80 micron wire position resolution with two planes after
correction for angle
o 200 microns resolution from tracking residuals
* studies with Fe55 source
o see effect of screening of wire
o confirmed with Garfield simulation
Beni
* FDC operational now, but not with normal gas mixture
o CO2 MFC died
o CDC running off FDC system with 50-50 mix
o 100% argon in CDC for a bit
* target retracted, upstream end of CDC accessible
o Beni: work on grounding, noise on CDC even after high frequency
filters on HV lines, likely noise from pre-amp cards
o sees slope on baseline
o need to talk to Tom if want to delay target re-insertion
* working with FSU student Brad Cannon on TOF calibration. Brad
on-site for two weeks
Bert
* 70% of all summer students at Lab working in Physics
o Rolf wants survey on student safety: labs, halls, storage buildings
o EH&S asking the questions
o look at hazard awareness, work control documents, required
training, knowledge of limits of work practices
o supervisors responsibility nominally
* HDList being used, please continue to do so
o see Bert if you need help
o posting a printed copy in work area a good practice
* about one week ago: golf cart event
o accelerator got stuck, hit bumper stops
o investigation on-going
o test brakes if using golf carts
Eugene, Solenoid
* planning test in two weeks
* test plan not firm
* an opportunity for test effect of field on detectors, be prepared if
this is a concern
* magnet on for 1 or 2 weeks
* considering re-doing cryostat on coil 4
o only one not done
o but coil 1 quenched first
Retrieved from
"https://halldweb.jlab.org/wiki/index.php?title=Private:Hall_D_Staff_Meeting,_July_27,_2015
<https://halldweb.jlab.org/wiki/index.php?title=Private:Hall_D_Staff_Meeting,_July_27,_2015&oldid=68916>"
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/halld-jlab/attachments/20150728/17413260/attachment.html>
More information about the Halld-jlab
mailing list